A 5.8 kg rock and a 2.20 x 10^-4 kg pebble are hel
Answer #1

c) F_rock = G*m_earth*m_rock/r_earth^2

d) F_rock = 6.67*10^-11*5.98*10^24*5.8/(6.38*10^6)^2

= 56.83 N


F_pebble = G*m_earth*m_pebble/r_earth^2

= 6.67*10^-11*5.98*10^24*2.2*10^-4/(6.38*10^6)^2

= 2.16*10^-3 N


e) a_rock = F_rock/m_rock

= (G*m_earth*m_rock/r_earth^2)/m_rock

= G*m_earth/r_earth^2

f) a_rock = 6.67*10^-11*5.98*10^24/(6.38*10^6)^2

= 9.8 m/s^2

a_pebble = G*m_earth/r_earth^2

= 6.67*10^-11*5.98*10^24/(6.38*10^6)^2

= 9.8 m/s^2